# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ IV-C.1 Least-Squares Estimation
Consider a narrowband BD-RIS-aided MISO system consisting of an $N$-antenna base station, an $M$-antenna BD-RIS working on reflecting mode, and a single-antenna user. In the uplink, denote the channels from the user to BD-RIS as $\mathbf{h}_{IR}\in\mathbb{C}^{M\times 1}$ and from the BD-RIS to the base station as $\mathbf{H}_{TI}\in\mathbb{C}^{N\times M}$. We assume the direct link from the user to the base station is blocked, $\mathbf{h}_{TR}=\mathbf{0}_{N\times 1}$, and focus only on the estimation of BD-RIS-aided channels. Then, the uplink channel from the user to the base station is given by
@@
Comparing $\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}$ and $\mathbf{H}_{\mathsf{cas}}^{\mathsf{D}}$, we could observe that the two cascaded channels have different dimensions and mathematical structures, such that the passive channel estimation designed for D-RIS does not work for BD-RIS. In the following, we will elaborate more on the difference in pattern design between BD-RIS and D-RIS. To do so, we first briefly revisit the channel estimation process [80], which generally works for any BD-RIS architectures and D-RIS.
@@
with $J\geq M^{2}$ to guarantee the successful estimation. This leads to the mean square error (MSE) as $\mathrm{err}_{\widehat{\mathbf{H}}_{\mathsf{cas}}^{\mathsf{BD}}}=\mathbb{E}\{\|\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}-\widehat{\mathbf{H}}_{\mathsf{cas}}^{\mathsf{BD}}\|_{\mathsf{F}}^{2}\}=\frac{N\sigma^{2}}{P_{\mathsf{u}}}\mathsf{tr}((\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}}(\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}})^{\mathsf{H}})^{-1})$, which indicates that the channel estimation performance depends purely on the design of $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}}$. We assume the BD-RIS has a non-reciprocal architecture, where the scattering matrices $\ddot{\mathbf{\Theta}}_{j}$, $\forall j\in\mathcal{J}$ are constrained by $\mathcal{T}_{\mathsf{non-recip}}$ in (70), for the ease of illustration and comparison. This motivates the formulation of the following problem
@@
where the training overhead is minimized as $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ without introducing estimation ambiguities. Similarly, for the case of D-RIS, we define the RIS pattern $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{D}}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{J}]\in\mathbb{C}^{M\times J}$, where $\bm{\theta}_{j}=[\ddot{\varTheta}_{j,1},\ldots,\ddot{\varTheta}_{j,M}]^{\mathsf{T}}$ collects the diagonal entries in $\ddot{\mathbf{\Theta}}_{j}=\mathsf{diag}(\ddot{\varTheta}_{j,1},\ldots,\ddot{\varTheta}_{j,M})$ of D-RIS at time slot $j$, with $|\ddot{\varTheta}_{j,m}|=1$, $\forall j\in\mathcal{J}$ when the RIS is lossless. To minimize the channel estimation error, the following problem is formulated
@@
where the training overhead is minimized as $J_{\mathsf{LS}}^{\mathsf{D}}=M$ here for the same reason as in problem (87). Problems (87) and (88) clearly show the difference between BD-RIS and D-RIS in pattern design. Specifically, for (88), one can directly use orthogonal matrices, such as the discrete Fourier matrix or the Hadamard matrix, to construct $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{D}}$ [15]. However, the solutions for (88) are not feasible for (87) due to the unique constraints of BD-RIS.
@@
[82, 83] $2M$ A two-phase estimation scheme by exploring the cascaded channel characteristics
[84] $3M$ A three-stage estimation scheme by exploring the property of non-diagonal scattering matrices
[85] Separate $M^{2}$ A decoupled estimation method based on Khatri-Rao Factorization
[86, 87] Much smaller than $M^{2}$ A decoupled estimation method based on Tucker decomposition and alternating least squares
[88] / A semi-blind estimation method that avoids the pilot-assisted training stage
@@
    [80, 81, 82, 83, 85, 86, 87, 88] focus on non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-recip}}$ and [84] focuses on non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-diag}}$.
@@ IV-C.2 Related Works and Discussions
To perfectly capture the constraints of BD-RIS architectures and facilitate the least-squares estimation, [80] has proposed a global optimal closed-form solution to (87), followed by a further extension to BD-RIS with hybrid/multi-sector modes in more general multi-user MIMO systems [81]. Based on the proposed pattern design, [80, 81] for the first time provide a comprehensive study for BD-RIS-aided systems from channel estimation to beamforming design and data transmission. However, the drawbacks of the method in [80, 81] are 1) the high training overhead, e.g., $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ for a non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-recip}}$, 2) the limited performance without exploring the structural characteristics of the cascaded channel, and 3) the channel estimation error for non-reciprocal BD-RIS that theoretically scales with the group size $\bar{M}$ [81]:
@@
which indicates that the estimation error increases with the circuit complexity of BD-RIS (i.e., a larger $\bar{M}$ leads to higher estimation error). To further reduce the channel estimation error, tensor decomposition has been applied in [85] based on the designed pattern in [80], which leads to the possibility of separate estimation of base station-RIS and RIS-user channels. Beyond that, two tensor decomposition based algorithms have been proposed in [86] with a milder requirement of the training overhead and enhanced estimation performance. [87] again explores the tensor decomposition for BD-RIS-aided channel estimation, which further facilitates the channel prediction with channel aging. [88] proposes a semi-blind channel and symbol estimation method that avoids the pilot-assisted estimation by exploring also the tensor decomposition for BD-RIS-aided channels. In addition, [82] has focused on the protocol design for a single-user MISO system and proposed a novel two-phase estimation scheme, where the cascaded channel related to the first element of BD-RIS is estimated, followed by fast estimation of the other channel coefficients. The two-phase estimation scheme explores the dependence between entries of the cascaded channel $\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}$ and thus helps to significantly reduce the training overhead from $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ to $J_{\mathsf{2-phase}}^{\mathsf{BD}}=2M$. This channel estimation scheme has been further extended to more general multi-user MIMO systems, with the training overhead in the same order as in D-RIS-aided system [83]. Besides, [84] has proposed a three-stage protocol for a non-reciprocal BD-RIS aided MISO system where the BD-RIS is characterized by a non-diagonal scattering matrix constrained by $\mathcal{T}_{\mathsf{non-diag}}$ in (66). Using the special signal flowing property in the non-diagonal scattering matrix, i.e., the signal impinging on one element is purely reflected by another one, the training overhead can be as low as $J_{\mathsf{3-stage}}^{\mathsf{ND}}=3M$. For clarity, the aforementioned works have been summarized in Table VI.
@@
The channel estimation study for BD-RIS-aided wireless systems is still at the very early stage. It is worth noting that, the passive channel estimation approach relies on the BD-RIS pattern design and results in a cascaded channel (or separate channels based on tensor decomposition whose process relies on BD-RIS constraints), both of which are related to BD-RIS architectures. In this sense, the existing literature [80, 81, 85, 86, 82, 83, 87, 84] based on non-reciprocal architectures cannot be readily extended to other BD-RIS architectures. Another critical issue for passive channel estimation is the training overhead that grows with the number of elements and the circuit complexity. Therefore, it remains unexplored, but it is important to develop BD-RIS pattern designs automatically adapting to architectures and to study better channel estimation schemes with affordable training overhead.
V BD-RIS with Hardware Impairments

### v2

@@ IV-C.1 Least-Squares Estimation
Consider a narrowband BD-RIS-aided MISO system consisting of an $N$-antenna base station, an $M$-antenna BD-RIS working on reflecting mode, and a single-antenna user. In the uplink, denote the channels from the user to BD-RIS as $\mathbf{h}_{IR}\in\mathbb{C}^{M\times 1}$ and from the BD-RIS to the base station as $\mathbf{H}_{TI}\in\mathbb{C}^{N\times M}$. We assume the direct link from the user to the base station is blocked [^5] In the case where a direct link exists, one can first turn OFF the BD-RIS by letting $\mathbf{\Theta}$ be a zero matrix and estimate the direct channel using well-established methods in conventional MIMO systems, and then estimate the cascaded channel related to BD-RIS by removing the contribution of the direct channel from the received data and using the proposed pattern design., $\mathbf{h}_{TR}=\mathbf{0}_{N\times 1}$, and focus only on the estimation of BD-RIS-aided channels. The uplink channel is thus
@@
Comparing $\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}$ and $\mathbf{H}_{\mathsf{cas}}^{\mathsf{D}}$, we could observe that the two cascaded channels have different dimensions and mathematical structures, such that the passive channel estimation designed for D-RIS does not work for BD-RIS. In the following, we will elaborate more on the difference in pattern design between BD-RIS and D-RIS. To do so, we first briefly revisit the channel estimation process [li2023channel], which generally works for any BD-RIS architectures and D-RIS.
@@
with $J\geq M^{2}$ to guarantee the successful estimation. This leads to the mean square error (MSE) as $\mathrm{err}_{\widehat{\mathbf{H}}_{\mathsf{cas}}^{\mathsf{BD}}}=\mathbb{E}\{\|\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}-\widehat{\mathbf{H}}_{\mathsf{cas}}^{\mathsf{BD}}\|_{\mathsf{F}}^{2}\}=\frac{N\sigma^{2}}{P_{\mathsf{u}}}\mathsf{tr}((\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}}(\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}})^{\mathsf{H}})^{-1})$, which indicates that the channel estimation performance depends purely on the design of $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{BD}}$. We assume the BD-RIS has a non-reciprocal architecture, where the scattering matrices $\ddot{\mathbf{\Theta}}_{j}$, $\forall j\in\mathcal{J}$ are constrained by $\mathcal{T}_{\mathsf{non-recip}}$ in (67), for the ease of illustration and comparison. This motivates the following problem
@@
where the training overhead is minimized as $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ without introducing estimation ambiguities. Similarly, for the case of D-RIS, we define the RIS pattern $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{D}}=[\bm{\theta}_{1},\ldots,\bm{\theta}_{J}]\in\mathbb{C}^{M\times J}$, where $\bm{\theta}_{j}=[\ddot{\varTheta}_{j,1},\ldots,\ddot{\varTheta}_{j,M}]^{\mathsf{T}}$ collects the diagonal entries in $\ddot{\mathbf{\Theta}}_{j}=\mathsf{diag}(\ddot{\varTheta}_{j,1},\ldots,\ddot{\varTheta}_{j,M})$ of D-RIS at time slot $j$, with $|\ddot{\varTheta}_{j,m}|=1$, $\forall j\in\mathcal{J}$ when the RIS is lossless. To minimize the channel estimation error, we have
@@
where the training overhead is minimized as $J_{\mathsf{LS}}^{\mathsf{D}}=M$ here for the same reason as in problem (81). Problems (81) and (82) clearly show the difference between BD-RIS and D-RIS in pattern design. Specifically, for (82), one can directly use orthogonal matrices, such as the discrete Fourier matrix or the Hadamard matrix, to construct $\mathbf{\Theta}_{\mathsf{all}}^{\mathsf{D}}$ [swindlehurst2022channel]. However, the solutions for (82) are not feasible for (81) due to the unique constraints of BD-RIS.
@@
[wang2024low, wang2025low] $2M$ A two-phase estimation scheme by exploring the cascaded channel characteristics
[samy2025low] $3M$ A three-stage estimation scheme by exploring the property of non-diagonal scattering matrices
[sokal2024decoupled] Separate $M^{2}$ A decoupled estimation method based on Khatri-Rao Factorization
[de2024channel, ginige2024efficient] Much smaller than $M^{2}$ A decoupled estimation method based on Tucker decomposition and alternating least squares
[de2024semi] / A semi-blind estimation method that avoids the pilot-assisted training stage
@@
    [li2023channel, li2024channel, wang2024low, wang2025low, sokal2024decoupled, de2024channel, ginige2024efficient, de2024semi] focus on non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-recip}}$ and [samy2025low] focuses on non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-diag}}$.
@@ IV-C.2 Related Works and Discussions
To perfectly capture the constraints of BD-RIS architectures and facilitate the least-squares estimation, [li2023channel] has proposed a global optimal closed-form solution to (81), followed by a further extension to BD-RIS with hybrid/multi-sector modes in more general multi-user MIMO systems [li2024channel]. Based on the proposed pattern design, [li2023channel, li2024channel] for the first time provide a comprehensive study for BD-RIS-aided systems from channel estimation to beamforming design and data transmission. However, the drawbacks of the method in [li2023channel, li2024channel] are 1) the high training overhead, e.g., $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ for a non-reciprocal BD-RIS constrained by $\mathcal{T}_{\mathsf{non-recip}}$, 2) the limited performance without exploring the structural characteristics of the cascaded channel, and 3) the channel estimation error for non-reciprocal BD-RIS that theoretically scales with the group size $\bar{M}$ [li2024channel]:
@@
which indicates that the estimation error increases with the circuit complexity of BD-RIS (i.e., a larger $\bar{M}$ leads to higher estimation error). To further reduce the channel estimation error, tensor decomposition has been applied in [sokal2024decoupled] based on the designed pattern in [li2023channel], which leads to the possibility of separate estimation of base station-RIS and RIS-user channels. Beyond that, two tensor decomposition based algorithms have been proposed in [de2024channel] with a milder requirement of the training overhead and enhanced estimation performance. [ginige2024efficient] again explores the tensor decomposition for BD-RIS-aided channel estimation, which further facilitates the channel prediction with channel aging. [de2024semi] proposes a semi-blind channel and symbol estimation method that avoids the pilot-assisted estimation by exploring also the tensor decomposition for BD-RIS-aided channels. In addition, [wang2024low] has focused on the protocol design for a single-user MISO system and proposed a novel two-phase estimation scheme, where the cascaded channel related to the first element of BD-RIS is estimated, followed by fast estimation of the other channel coefficients. The two-phase estimation scheme explores the dependence between entries of the cascaded channel $\mathbf{H}_{\mathsf{cas}}^{\mathsf{BD}}$ and thus helps to significantly reduce the training overhead from $J_{\mathsf{LS}}^{\mathsf{BD}}=M^{2}$ to $J_{\mathsf{2-phase}}^{\mathsf{BD}}=2M$. This channel estimation scheme has been further extended to more general multi-user MIMO systems, with the training overhead in the same order as in D-RIS-aided system [wang2025low]. Besides, [samy2025low] has proposed a three-stage protocol for a non-reciprocal BD-RIS aided MISO system where the BD-RIS is characterized by a non-diagonal scattering matrix constrained by $\mathcal{T}_{\mathsf{non-diag}}$ in (63). Using the special signal flowing property in the non-diagonal scattering matrix, i.e., the signal impinging on one element is purely reflected by another one, the training overhead can be as low as $J_{\mathsf{3-stage}}^{\mathsf{ND}}=3M$. For clarity, the aforementioned works have been summarized in Table VI.
@@
The channel estimation study for BD-RIS-aided wireless systems is still at the very early stage. It is worth noting that, the passive channel estimation approach relies on the BD-RIS pattern design and results in a cascaded channel (or separate channels based on tensor decomposition whose process relies on BD-RIS constraints), both of which are related to BD-RIS architectures. In this sense, the existing literature [li2023channel, li2024channel, sokal2024decoupled, de2024channel, wang2024low, wang2025low, ginige2024efficient, samy2025low] based on non-reciprocal architectures cannot be readily extended to other BD-RIS architectures. Moreover, in practice, BD-RIS inevitably has hardware impairments that will induce new mathematical constraints and communications models, as will be detailed in Section VI. Therefore, those works based on BD-RIS with perfect hardware may not be applicable to BD-RIS with practical hardware impairments. Another critical issue for passive channel estimation is the training overhead that grows with the number of elements and the circuit complexity. Therefore, it remains unexplored, but it is important to develop BD-RIS pattern designs automatically adapting to architectures and hardware impairments, and to study better channel estimation schemes with affordable training overhead.
V Benefits of BD-RIS
Thanks to the flexible interconnections between elements to support various architectures and modes, BD-RIS has multiple benefits, such as enhancing channel gain, increasing transmission quality, enlarging coverage, etc. In this section, we summarize the key benefits of BD-RIS, each of which is supported by numerical results.
V-A Boosting Received Power and Rates
Compared to (lossless) D-RIS which can only manipulate the phase shift of the diagonal entries in the scattering matrix, BD-RIS has higher flexibility in manipulating both amplitude and phase shift of diagonal and off-diagonal entries in its scattering matrix. This flexibility boosts the performance of various systems, such as increasing the received power by up to 62% for SISO as shown in Fig. 4 and increasing the sum-rate for MU-MISO by up to 43% as shown in Fig. 16.
Figure 16: Sum-rate of a 4-user system with the aid of a 32-element BD-RIS versus transmit power for different BD-RIS architectures. Channels through BD-RIS follow Rician fading with a Rician factor 2 dB. The direct transmitter-user links are assumed to be blocked. The transmitter-RIS distance is set as 50 m and the RIS-user distance is set as 2.5 m [wu2024optimization].
V-B Enabling Low-Complexity Architectures with High Performance
The flexibility provided by the interconnections between BD-RIS elements enables various circuit topology designs to support different architectures as illustrated in Sections III-D and III-E. More importantly, this flexibility enables architectures with least hardware complexity (i.e., number of tunable components) to achieve optimal performance in various systems.
For the MISO system, [nerini2024beyond] has theoretically proven that the least circuit complexity for MISO optimal BD-RIS is
$C_{\mathsf{MISO}}^{\mathsf{opt}}=2M-1.$ (84)
This condition essentially forms a tree-connected architecture, and two representative examples are arrowhead BD-RIS and tridiagonal BD-RIS. Based on this conclusion, [nerini2023pareto] has analyzed the Pareto frontier between performance and circuit complexity for a SISO system. Fig. 17 shows that tree-connected architecture can reach the performance achieved by fully-connected BD-RIS with the lowest possible complexity, thus achieving the best performance-complexity trade-off.
Figure 17: Pareto frontier between performance and circuit complexity including single-, group-, fully-, forest-, and tree-connected BD-RIS, with different group sizes $\bar{M}$ labeled on two sides of the curve [nerini2023pareto]. The direct transmitter-receiver channel is assumed to be blocked and channels through BD-RIS follow i.i.d. Rayleigh fading.
Figure 18: Pareto frontier between performance and circuit complexity including single-, group-, fully-, stem-, and band-connected BD-RIS, with different group sizes $\bar{M}$ and stem/band width $q$ labeled on two sides of the curves. The direct transmitter-user channels are assumed to be blocked and channels through BD-RIS follow Rician fading with a Rician factor 2 dB. The transmitter-RIS distance is set as 50 m and the RIS-user distance is set as 2.5 m. ($N=K=4$, $M=64$, $P=10$ dBm) [wu2025beyond].
For more general multi-user MIMO systems, [wu2025beyond] has theoretically proved that the least circuit complexity for multi-user MIMO optimal BD-RIS is
$C_{\mathsf{MU-MIMO}}^{\mathsf{opt}}=\min\left\{D,\frac{M}{2}\right\}\left(2M-2\min\left\{D,\frac{M}{2}\right\}+1\right),$ (85)
where $D=\min\{\sum_{k\in\mathcal{K}}N_{k},N\}$ is the DoF of the multi-user MIMO channel. Two representative examples which can reach such condition are band- and stem-connected BD-RIS with band and stem width $q=2\min\{D,\frac{M}{2}\}-1$. Fig. 18 reports the Pareto frontier between performance and circuit complexity achieved by BD-RIS-aided multi-user MISO systems, indicating that band/stem-connected BD-RIS with proper band/stem width $q$ can reach the performance achieved by fully-connected BD-RIS, with much less circuit complexity.
Figure 19: Sum-rate for a BD-RIS-aided multi-user MISO system versus the number of elements $M$. “CW-FC”, “CW-GC”, and “CW-SC”, respectively, refer to cell-wise fully-, group-, and single-connected. Channels through BD-RIS follow Rician fading with a Rician factor 0 dB. The transmitter-user channels are fully blocked. The transmitter-RIS distance is set as 100 m and the RIS-user distance is set as 10 m. The group size for hybrid and multi-sector modes with CW-GC architectures are respectively set as 4 and 8. The number of sectors for the multi-sector mode is set as 4. $N=K=4$ with four users distributed evenly across the sectors covered by BD-RIS, that is, all users on the same side for reflecting mode, every two users on one side for hybrid mode, and one user within each sector for multi-sector mode.
V-C Enabling Flexible Modes with Highly-Directional Wireless Coverage
The inter-element connections in BD-RIS not only boost the channel strength and system performance, but also provide possibility to flexibly arrange the locations and orientations of elements. This leads to different modes with enlarged wireless coverage, such as hybrid mode proposed in [li2022beyond] and multi-sector modes proposed in [li2023beyond]. Results in Fig. 19 demonstrate that a reflecting BD-RIS can achieve up to 75% of sum-rate improvement over D-RIS (single-conn.) when $M=128$. More importantly, a 4-sector BD-RIS with cell-wise fully-connected architecture can increase the sum-rate by 150% over STAR-RIS with $M=128$, thanks to the higher antenna gains provided by more directional antennas.
V-D Providing Orders of Magnitude Gains in Distributed Deployments
In most existing literature, an RIS has been regarded as an antenna array whose inter-element spacing is comparable with (typically smaller than) the wavelength. This setting results in a localized RIS whose elements are localized in a specific site. The localized RIS has been widely studied in far-field scenarios, where the wireless channels related to each RIS element share approximately the same large-scale fading, as shown in Fig. 20(a). In this sense, the gain of BD-RIS over D-RIS mainly comes from the better exploration of small-scale fading effects. To understand if the joint effects of small- and large-scale fading in wireless channels can provide additional flexibility to be captured by BD-RIS architectures, [nerini2025localized] has proposed the concept of distributed RIS. In the distributed RIS, elements are distributed over a wide region with inter-element spacing much larger than the wavelength, such that the large-scale fading between transmitter/receiver and each RIS element can be quite different, as shown in Fig. 20(b).
To quantify the benefit of distributed RIS, [nerini2025localized] has derived the gain of distributed over localized fully-connected BD-RIS in SISO systems under Rayleigh fading channels, which is bounded by
where $d_{R}$ and $d_{T}$, respectively, denote the distance between the localized RIS and the receiver/transmitter; $\mathbf{d}_{R}\in\mathbb{R}^{M\times 1}$ and $\mathbf{d}_{T}\in\mathbb{R}^{M\times 1}$, respectively, denote the distance between all $M$ elements of the distributed RIS and the receiver/transmitter; $a$ denotes the path-loss exponent. The bounds in (86) indicate that to guarantee $G^{\mathsf{dis}}>1$, a sufficient condition is to have $d_{R}d_{T}>\sqrt[a]{M^{2}}\min\{\mathbf{d}_{R}\}\min\{\mathbf{d}_{T}\}$ and a necessary condition is to have $d_{R}d_{T}>\min\{\mathbf{d}_{R}\}\min\{\mathbf{d}_{T}\}$, both of which are mild and can be easily achieved. This is also numerically evaluated in Fig. 21(a), showing that $G^{\mathsf{dis}}$ grows exponentially with $a$ and reaches as high as several orders of magnitude. The impact of locations of the receiver is reported in Fig. 21(b), showing that the distributed arrangement has the most significant performance benefit when the receiver is far from localized BD-RIS while being close to distributed BD-RIS.
(a) Gain of distributed over localized BD-RIS $G^{\mathsf{dis}}$
(b) Gain of distributed over localized BD-RIS $G^{\mathsf{dis}}$
Figure 21: $G^{\mathsf{dis}}$ (in dB) for different values of path-loss exponent, number of RIS elements, and locations of receiver [nerini2025localized].
V-E Enabling Simultaneously Optimal Transmissions for Uplink and Downlink with Non-Reciprocal Architectures
One key property of non-reciprocal BD-RIS is that its admittance/impedance/scattering matrices are asymmetric, which naturally break the uplink-downlink reciprocity of the wireless channels. For example, given an uplink BD-RIS-aided MISO channel $\mathbf{h}_{\mathsf{up}}=\mathbf{H}_{TI}\mathbf{\Theta}\mathbf{h}_{IR}$ as shown in (75), the downlink channel $\mathbf{h}_{\mathsf{down}}=\mathbf{h}_{RI}\mathbf{\Theta}\mathbf{H}_{IT}=\mathbf{h}_{IR}^{\mathsf{T}}\mathbf{\Theta}\mathbf{H}_{TI}^{\mathsf{T}}$ is not equal to the transpose of the uplink channel, i.e., $\mathbf{h}_{\mathsf{down}}\neq\mathbf{h}_{\mathsf{up}}^{\mathsf{T}}$ since $\mathbf{\Theta}\neq\mathbf{\Theta}^{\mathsf{T}}$. This property provides unique benefits to scenarios where uplink and downlink transmissions behave differently, such as full-duplex systems [zhang2016full].
To visualize the benefit of applying non-reciprocal BD-RIS in full-duplex systems, [li2024non] has derived the general RIS-aided full-duplex system model, and the theoretical conditions for non-reciprocal BD-RIS to simultaneously maximize the received powers of the signal of interest in the uplink and downlink. Results in the top two figures of Fig. 22 show that, when the uplink and downlink users are not aligned, both the impinging and reflected beams of non-reciprocal BD-RIS can exactly point to the directions of the signal of interest with the maximum power, while those of reciprocal BD-RIS fail to point to the downlink user. Meanwhile, the bottom two figures in Fig. 22 show that, only when the uplink and downlink users are aligned, both the impinging and reflected beams of reciprocal and non-reciprocal BD-RIS can exactly point to the directions of signal of interest. Due to the unique property of non-reciprocal BD-RIS, if a wave hits the non-reciprocal BD-RIS from one direction, the surface behaves differently than if it hits from the opposite direction. This finally enables an uplink user and a downlink user at different locations to optimally communicate with the same full-duplex base station. While [li2024non] primarily focuses on a single-antenna scenario, the benefit of non-reciprocal BD-RIS in full-duplex systems has been recently shown in more general multi-user multi-antenna scenarios [liu2024non].
Figure 22: The impinging and reflected beam patterns of reciprocal and non-reciprocal BD-RISs. The channels between devices are assumed to be pure LoS. The base station (marked as “BS”) is located at $\frac{\pi}{6}$; the downlink user is located at $\frac{\pi}{2}$; and the uplink user is located at $\frac{2\pi}{3}$ in the top figures and at $\frac{\pi}{2}$ in the bottom figures [li2024non].
V-F Providing Enhanced Gains in Dual-Polarized Systems
Existing BD-RIS literature has focused on uni-polarized systems for the ease of analysis. However, modern MIMO systems use dual-polarized antenna arrays to have more antennas within limited space to provide more beamforming gains [kim2010limited]. Dual-polarized D-RIS-aided systems have been studied in [han2021dual, zheng2024ris], while it remains unknown if BD-RIS still has performance gains over D-RIS in dual-polarized systems. Taking into account this practical consideration, [nerini2025dual] has modeled and analyzed a dual-polorized BD-RIS-aided SISO system, where the BD-RIS has half vertically polarized elements and half horizontally polarized elements, and the receiving and transmitting antennas can have the same or opposite polarizations, as illustrated in Fig. 23.
Figure 24: Gain of BD-RIS over D-RIS $\lim_{M\rightarrow\infty}G^{\mathsf{dual-polarized}}$ as a function of $\chi$ when the transmitter and receiver have the same/opposite polarization and with Rayleigh/LoS channels [nerini2025dual].
When the transmitter and receiver have the same polarizations, as illustrated in Fig. 23(a), the limit of the power gain of dual-polarized fully-connected BD-RIS over D-RIS remains the same as the uni-polarized case [shen2021]. That is, dual-polarized BD-RIS can increase the received power in SISO by up to 62% compared to dual-polarized D-RIS. When the transmitter and receiver have opposite polarizations, the limit of the power gain of dual-polarized fully-connected BD-RIS over D-RIS varies with channel fadings, and is expressed as [nerini2025dual]
where $0<\chi<1$ denotes the inverse of the cross-polar discrimination. This analytical result indicates that BD-RIS can offer a gain $G^{\mathsf{dual-polarized}}>1$ for both Rayleigh fading and LoS channels. To visualize the impact of $\chi$ on performance gain, Fig. 24 illustrates the gain of BD-RIS over D-RIS as a function of $\chi$. These analytical results show that BD-RIS has more significant benefits over D-RIS with a smaller $\chi$, as small values of $\chi$ increase disparity between the channel entries.
VI BD-RIS with Hardware Impairments